the new drug was administered to 15 patients, the following

drops in systolic blood pressure were registered: 12, 8, 15, 9, 10, 11,

14, 13, 6, 9, 14, 12, 18, 14 and 16 points. Assuming that the drops in

systolic blood pressure are normally distributed.

(a) Construct a 99% confidence interval for the mean drops in blood

pressure.

(b) The manufacturer of a new drug claims that it will lower blood

pressure by 13 points on the average. Is the claim sustained at

the 0.01 level of significance?



1
Expert's answer
2021-01-29T04:57:37-0500
"12, 8, 15, 9, 10, 11,\n\n14, 13, 6, 9, 14, 12, 18, 14,16"

"\sum x_i=181, n=15"

"\bar{x}=\dfrac{181}{15}\approx12.067"

"\sum x_i^2=2333"

"s^2=\dfrac{1}{n-1}\big(\sum x_i^2-\dfrac{1}{n}(\sum x_i)^2\big)"

"=\dfrac{1}{15-1}\big(2333-\dfrac{1}{15}(181)^2\big)\approx10.6381"

"s=\sqrt{s^2}=\sqrt{10.6381}\approx3.2616"

(a) The provided sample mean is "\bar{x}=12.067" and the sample standard deviation is "s=3.2616." The size of the sample is "n=15" and the required confidence level is "99\%."

The number of degrees of freedom are "df=n-1=15-1=14," and the significance level is "\alpha=0.01." Based on the provided information, the critical t-value for "\alpha=0.01" and "df=14" degrees of freedom is "t_c=2.9768."

The 99% confidence for the population mean "\mu" is computed using the following expression


"CI=\big(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}}\big)"

"=\big(12.067-2.9768\times\dfrac{3.2616}{\sqrt{15}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}}\big)"

"=(9.560, 14.574)"



(b) The provided sample mean is "\bar{x}=12.067" and the sample standard deviation is "s=3.2616," and the size of the sample is "n=15."

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq13"

"H_1:\mu<13"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

The number of degrees of freedom are "df=n-1=15-1=14," and the significance level is "\alpha=0.01." Based on the provided information, the critical t-value for "\alpha=0.01" and "df=14" degrees of freedom is "t_c=-2.624."

The rejection region for this left-tailed test is "R=\{t:t<-2.624\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{12.067-13}{3.2616/\sqrt{15}}=-1.1082"

Since it is observed that "t=-1.1082>-2.624=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 13, at the 0.01 significance level.

Use the P-value approach.

The P-value "t=-1.1082, df=14, \alpha=0.01," left-tailed, is "p=0.1433." Since "p=0.1433>0.01=\alpha," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 13, at the 0.01 significance level.

The claim is sustained at the 0.01 level of significance.